OpenStudy (anonymous):

Use Newton's Method to approximate the root of x lnx=2 accurate to five decimal places. Begin by sketching a graph. this one i am having problems with

OpenStudy (anonymous):

\[x\ln(x)=2\] \[x\ln(x)-2=0\] make a guess

OpenStudy (anonymous):

doesnt even have to be that good of a guess. i guess 2 and get \[f(2)=2\ln(2)-2=-.61370...\]

OpenStudy (anonymous):

ohhh okay i understand what to do now...i can pick any number right?

OpenStudy (anonymous):

the derivative of \[x\ln(x)-2\] is \[1+\frac{1}{x}\]

OpenStudy (anonymous):

yeah some number that will give you something close to 0. doesn't even have to be that close. i picked 2. next number will be \[2-\frac{f(2)}{f'(2)}\]

OpenStudy (anonymous):

so i get \[2-\frac{-.6137}{(1+\frac{1}{-.6137})}\]

OpenStudy (anonymous):

now i need a calculator for sure

OpenStudy (anonymous):

gives 1.02504

OpenStudy (anonymous):

next number is \[1.02504-\frac{f(1.02504)}{(1+\frac{1}{1.02504})}\]

OpenStudy (anonymous):

oh crap! this is a tricky one

OpenStudy (anonymous):

it is bt i think they want an approximation that is close to 2

OpenStudy (anonymous):

if you look at the graph you will see that the tangent lines head off in the wrong direction. you need to pick somehing bigger than 2, say 2.5. that should work. newton's method does not always work right and if you look at the graph you will see why. tangent lines get farther from the roots, not close. try looking here http://www.wolframalpha.com/input/?i=y%3Dx*ln%28x%29-2